OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? x − y = 8 −x − y = −7 8 − y −8 − y 8 + y −8 + y

OpenStudy (cwrw238):

x - y = 8 add y to both sides

OpenStudy (anonymous):

isolating the x, means moving all other constants and variables to the other side of the equation. This will leave you with x=... They are asking what x= for the first equation, which you substitute into the second equation.

OpenStudy (anonymous):

oh

OpenStudy (cwrw238):

x - y + y = 8 + y can you simplify this

OpenStudy (anonymous):

yes

OpenStudy (cwrw238):

part of it cancels right?

OpenStudy (cwrw238):

so which one of the choices is it?

OpenStudy (anonymous):

8 + y

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

thank you

OpenStudy (cwrw238):

yw